OpenStudy (ksaimouli):

find the tangent y=tan^-1 x at the point (1,pi/4)

OpenStudy (anonymous):

\[\tan(y)=x\] \[\frac{dx}{dy}=\frac{1}{\cos^2(y)}\] Are you fine with differentiating tan?

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):